OpenStudy (anonymous):

PLEASE HELP!!! find the equation of the line in slope-intercept form that is perpendicular to y=-1/2x+6 and passes through (3,-1)

jimthompson5910 (jim_thompson5910):

What's the slope of y=-1/2x+6 ?

jimthompson5910 (jim_thompson5910):

If you're not sure, remember the slope intercept form y = mx+b m is the slope

OpenStudy (anonymous):

The slope is -1/2

jimthompson5910 (jim_thompson5910):

flip the fraction: -1/2 -----> -2/1 = -2 Then flip the sign: -2 -----> +2 = 2

jimthompson5910 (jim_thompson5910):

The original slope is -1/2 The perpendicular slope is 2

jimthompson5910 (jim_thompson5910):

So we want an equation that A) has a slope of 2 B) goes through the point (3,-1)

jimthompson5910 (jim_thompson5910):

do you know where to go from here?

OpenStudy (anonymous):

no I don't get it

jimthompson5910 (jim_thompson5910):

do you see how I got the slope of 2?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

now we use y = mx+b to find the equation

OpenStudy (anonymous):

so now it would be y=2x+b?

jimthompson5910 (jim_thompson5910):

We have a slope m = 2 We have a point (x,y) = (3,-1) Use them both to plug them into y = mx+b to find the equation y = mx+b y = 2x+b ... plug in the slope m = 2 -1 = 2*(3) + b ... plug in the point (x,y) = (3,-1) b = ???

jimthompson5910 (jim_thompson5910):

yes you're on the right track

OpenStudy (anonymous):

does B =-7?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

your equation is therefore y = 2x - 7

OpenStudy (anonymous):

Thank you!!!

jimthompson5910 (jim_thompson5910):

you're welcome